A project was budgeted at $106915. The project is now being executed. The following figures have been determined:

PV:$36282

EV:$62637

AC:$50020

Assume that the cost variance was caused by the purchase of computer hardware last month, but that you do not expect any more variances from your original spending plan going forward. To the nearest dollar, what estimate at completion (EAC) do you predict?
Answer #1

EAC or estimate At completion can be calculated by dividing budget at completion by CPI

EAC = BAC/CPI

Budget at Completion = $ 106915

CPI = EV/AC = 62637/50020 = 1.252239

So, EAC = 106915/1.252239 = $85379.06 ~$85379 is the estimate at completion.